Among the 250 employees of the local office of an international insurance company, 182 are whites, 51 are blacks, and 17 are Orientals. If we use proportional allocation to select a stratified random grievance committee of 15 employees, how many employees must we take from each race?
1
Expert's answer
2020-10-01T15:41:08-0400

Given :

Total : 250

number of Blacks : 51

number of Whites 182

Number of Orientals: 17

Solution:


proportional allocation to select a stratified random grievance committee of 15 employees of Blacks= 51*"\frac{15}{250}= 3.06" ~3


proportional allocation to select a stratified random grievance committee of 15 employees of Whites= 182*"\frac{15}{250}=10.92" ~11


proportional allocation to select a stratified random grievance committee of 15 employees of Orientals: 17*"\frac{15}{250}=1.02" ~1